OpenStudy (anonymous):

find the eqn of a plane that passes through the point (1,2,3) and is parallel to x-y+z=100

OpenStudy (anonymous):

if it is parallel, then it will be in the form of x-y+z=c where c is a constant. since it has to pass through (1,2,3) we can find c by plugging (1,2,3) into the equation i.e. x=1,y=2,z=3. so (1-2+3=2) so the equation is x-y+z=2

OpenStudy (anonymous):

ok thank u :) thats what i got i was just making sure x-y+z-2=0

OpenStudy (anonymous):

no problem, what kind of math are you doing? i am learning this stuff at the moment from an online course

OpenStudy (anonymous):

calc 3

OpenStudy (anonymous):

do u know anything about curvatures?